### v1

@@ Abstract
Since Adam was introduced, several novel adaptive optimizers for deep learning have been proposed. These optimizers typically excel in some tasks but may not outperform Adam uniformly across all tasks. In this work, we introduce Meta-Adaptive Optimizers (MADA), a unified optimizer framework that can generalize several known optimizers and dynamically learn the most suitable one during training. The key idea in MADA is to parameterize the space of optimizers and search through it using hyper-gradient descent. Numerical results suggest that MADA is robust against sub-optimally tuned hyper-parameters, and outperforms Adam, Lion, and Adan with their default hyper-parameters, often even with optimized hyper-parameters. We also propose AVGrad, a variant of AMSGrad where the maximum operator is replaced with averaging, and observe that it performs better within MADA. Finally, we provide a convergence analysis to show that interpolation of optimizers (specifically, AVGrad and Adam) can improve their error bounds (up to constants), hinting at an advantage for meta-optimizers.
@@ 1 Introduction
The choice of an optimization algorithm plays a critical role in determining the downstream performance of a machine learning model. Adaptive moment optimizers are the most preferred class of optimizers employed in most learning tasks such as training Large Language Models (LLM) (Brown et al., 2020; Touvron et al., 2023) and Diffusion Models (Rombach et al., 2022). In particular, Adam (Kingma and Ba, 2015) is still the “go-to” optimizer in LLM training, despite the emergence of many other optimizers since then.
@@
In this work, we introduce the concept of a parameterized optimizer, which can be viewed as a unified parameterization of a collection of given optimizers. Specifically, the updating rules of a parameterized optimizer are a convex combination of those of the underlying optimizers, where the coefficient assigned to each underlying optimizer is learnable. To make this concept operational, we propose the meta-adaptive optimizer (MADA), which combines the parameterized optimizer with hyper-gradient descent (Baydin et al., 2018) to learn the optimizer coefficients. MADA dynamically adjusts the parameterized optimizer coefficients during training, and effectively adapts the optimizer choice to the learning task. While MADA bears some connections to optimizer search methods, such as (Chen et al., 2023), it does not solely output a final optimizer state. It dynamically adapts the optimizer to the current neighborhood of the loss landscape on-the-fly, removing the need for an offline optimizer search stage, or an outer hyper-parameter selection loop.
Contributions. In this paper, we make the following contributions:
  • We introduce the concept of parameterized optimizer, which takes a collection of existing optimizers, and unifies them into a single optimizer, combining the individual update rules through learnable coefficients.
@@
  • We propose a new optimizer, AVGrad, modifying AMSGrad (Reddi et al., 2018) by replacing the maximum operator on the second-order moments with time-averaging, which results in better performance within the MADA framework. Moreover, we provide a convergence analysis for AVGrad as well as its interpolations with Adam, demonstrating that optimizer interpolations can improve error bounds up to constant factors, providing evidence in favor of parameterized optimizers. We further show that MADA converges to AVGrad in a simple example where it is known to perform better than Adam, providing evidence for the effectiveness of MADA in adapting the optimizer to the task (see Appendix C).
@@
  • We develop a specific parameterized optimizer, which interpolates between Adam (Kingma and Ba, 2015), AVGrad, Yogi (Zaheer et al., 2018), Adan (Xie et al., 2023), and Lion (Chen et al., 2023). We empirically compare MADA, which is based on this parameterized optimizer, against Adam, Lion, Adan, and HyperAdam [^1] Throughout this paper we will refer to the version of Adam that uses hyper-gradients to tune $\beta_{1}$ and $\beta_{2}$ parameters, as in Chandra et al. (2022), as HyperAdam., and show that MADA generally outperforms all baselines. We also illustrate the robustness of MADA to initial hyper-parameters and analyze the evolution of hyper-parameters during training.
@@
Related work. Motivated by the high cost of training large language models, a large number of novel optimizers have been proposed in recent years to speed up the training, increase generalization performance or train more resource-efficient models. As mentioned before, Adam (Kingma and Ba, 2015) is the most commonly employed optimizer, and succeeding methods, in general, try to improve upon it under different scenarios. Reddi et al. (2018); Zaheer et al. (2018) propose AMSGrad and YOGI, respectively, to fix Adam’s potentially non-decreasing effective learning rate. Xie et al. (2023); Dozat (2016) introduce Adan and Nadam to replace heavy-ball momentum in Adam with Nesterov momentum. You et al. (2017, 2020) propose LARS and LAMB to improve the performance of Adam in large-batch regime. Heo et al. (2021) proposes AdamP to avoid premature decay of scale-invariant weights. AdaBound (Luo et al., 2019) and AdaBelief (Zhuang et al., 2020) tries to estimate a more stable second-order moment term. Foret et al. (2021) proposes sharpness-aware minimization (SAM) and Chen et al. (2020) proposes Padam to increase the generalization performance of the trained models. Liu et al. (2020) stabilizes the training by reducing gradient variance. The work in (Chen et al., 2023) is similar in spirit to our work, in that it introduces a method to symbolically search a space of optimizers; however unlike our work they consider an offline search method, whereas our method learns the optimizer during actual model training, and uses hyper-gradients.
@@ 3 Meta-Adaptive Optimizers
4 AVGrad and Convergence Analysis
Reddi et al. (2018) noted convergence issues in exponential moving average algorithms such as Adam. In particular, there were simple convex settings where Adam fails to converge to the optimal solution because it could lead to effectively increasing learning rates; to prevent this they introduced AMSGrad. However, while incorporating AMSGrad into MADA, we have observed a decrease in validation performance. We conjecture that this is caused by the maximum operator in the second-moment term. Specifically, note that back-propagation of gradients through $\max(a,b)$ corresponds to a simple routing operation to the larger one of $a$ and $b$, where the smaller one is passed 00 gradients. In AMSGrad, we have $v_{t}=\max\left\{v_{t-1},\bar{v}_{t}\right\}$ which means that for most steps the first term will be greater, causing insufficient hyper-gradient updates on $\beta_{2}$ parameter through $\bar{v}_{t}$ path [^4] To be accurate, considering the example parameterization (3), $\tilde{v}_{t}$ term provides another path for hyper-gradients on $\beta_{2}$. However in practice we have observed that when AMSGrad is used, $\rho$ parameter tends to vanish, diminishing the effect of this path as well..
To remedy this, we replaced the maximum operator with time-averaging in AMSGrad (calling the resulting algorithm AVGrad), and observed that it leads to a better performance in terms of the validation loss. The idea of using averaging first originated in Reddi et al. (2018); however, to the best of our knowledge, this algorithm was not analyzed before, theoretically or empirically.
Due to the complexity of our overall parameterization, deriving theoretical guarantees for MADA at its full scope is challenging. Nevertheless, in this section, we are able to examine the theoretical properties of interpolations between AVGrad and Adam. Specifically, we show that the interpolated optimizers can improve the convergence rate of either optimizer up to constant factors, demonstrating the potential value of the parameterized optimizer formulation also in the theoretical front.
4.1 AVGrad and the interpolated optimizer
@@
where $\tilde{v}_{t}$ is the running average of past $\overline{v}_{t}$’s. We provide the convergence analysis of AVGrad in Appendix B.
In what follows, we focus on the interpolation between Adam and AVGrad, with the interpolation coefficient $\rho_{t}$. Specifically, we replace last line (5) above with $v_{t}:=\rho_{t}\bar{v}_{t}+(1-\rho_{t})\tilde{v}_{t}$ where $\rho_{t}$ interpolates between second-order moments of Adam and AVGrad.
@@
In Appendix C, we use MADA to solve the simple convex problem given in (Reddi et al., 2018), which is an example where Adam fails. We found that MADA quickly converges to AVGrad, and thus to the optimum, by adapting $\rho_{t}\rightarrow 0$, even when we initialize it from Adam ($\rho_{0}=1$).
@@ 4.2 Convergence of the interpolated optimizer
In this subsection, we present the convergence bound for the interpolation of AVGrad and Adam (we defer the exact details to Appendix A). We make the following standard assumptions: $F$ is bounded from below, is $L$-smooth, and with stochastic gradients that are bounded almost surely.
Défossez et al. (2022) proposed a unified analysis for the convergence analysis of Adagrad and Adam in the non-convex setting. Since AVGrad can be seen as a version of Adagrad, the proof technique of Défossez et al. (2022) can be adapted to AVGrad with some modifications. Here we state the convergence result of the interpolated optimizer without momentum, that is when $\beta_{1}=0$. The extension to the case with momentum can be done in the same way we extend Theorem 2 to Theorem 3, similar to (Défossez et al., 2022).
@@
We provide the precise statement in Appendix A.
Remark. Aligned with the condition in Proposition 1, the contribution from $v_{t}$ in Theorem 1 is multiplied with $1/t$ in order to avoid divergent terms in the bound. Observe that when $\rho=0$ or $\rho=1$ we recover the convergence rates for AVGrad (see Theorem 2 in Appendix B) and Adam (Défossez et al., 2022) respectively. We note that for the right-hand side to remain bounded, one would need either $0\leq\rho\leq\beta_{2}$ (all terms vanish as in AVGrad), or $\rho=1$ (a constant term remains as in Adam). To gain more insights, in Figure 2 and 2, we plot the value of the bound in Theorem 1 with respect to $\rho$ for two representative cases (see Appendix A for more details). The first case with larger $T$ represents a more favorable setting for AVGrad due to the non-vanishing term in Adam bound, while the second is more favorable for Adam due to its $O(1/T)$ rate in the vanishing terms. In both cases, the interpolated optimizer provides the best convergence bound when $\rho=\beta_{2}$, which suggests an advantage for the interpolated optimizers.
@@ 5 Experiments
In our experiments, we aim to answer how the generalization and downstream performance of MADA compares against fixed optimizers, and how robust MADA is to poor hyper-parameter initializations compared to fixed optimizers. We focus on the pre-training setting training models from scratch. We also try to gain insights into the behavior of MADA by monitoring the evolution of the optimizer coefficients. Additional details of our experimental setup, and results can be found in Appendix C.
Figure 3: Validation loss on OpenWebText for GPT-2 (125M) model.
@@
Figure 3: Validation loss on OpenWebText for GPT-2 (125M) model.
5.1 A concrete parameterized optimizer
5.2 Experimental setting
Data and Model. In the recent years, auto-regressive models have been widely used for benchmarking and algorithm evaluation (Radford et al., 2019; Liu et al., 2023). Motivated by this, we evaluate MADA on the causal language modeling task with GPT-2, over two datasets: Shakespeare (Karpathy, 2015), and OpenWebText (Gokaslan and Cohen, 2019). On the Shakespeare dataset, we train a 6 layer transformer with context size of 256 (11M parameters). On OpenWebText, we train GPT-2 (small) with 12 layers, 1024 context size (125M parameters) and GPT-2 (medium) with 24 layers, 1024 context size (335M parameters). We use nanoGPT (https://github.com/karpathy/nanoGPT) codebase for the implementation.
Baselines and learned optimizers. On OpenWebText, we compare MADA to Adam (Kingma and Ba, 2015), HyperAdam (Chandra et al., 2022) (Adam with $\beta_{1}$ and $\beta_{2}$ parameters learned through hyper-gradients), Adan (Xie et al., 2023), Lion (Chen et al., 2023), and AVGrad. We keep the learning hyper-parameters such as learning rate schedule and weight decay identical, except for Lion where we choose a lower initial learning rate as suggested in (Liu et al., 2023). For a fair comparison, for all optimizers, we use the same codebase based on Chandra et al. (2022). For OpenWebText experiments, we use established parameters for Adam ($\beta_{1}=0.9,\beta_{2}=0.95,\epsilon=10^{-6}$) and also use these values as the initial parameters for MADA, HyperAdam, and AVGrad; for other methods we use the parameters suggested in respective papers; and measure the validation loss. For Shakespeare experiments, we compare MADA to Adam and HyperAdam; the relatively small model size in this experiment allows us to sweep the grid of initial $\beta$ parameters and compare the final training loss across many different hyper-parameter choices. In some experiments, we also evaluate the final state of MADA statically, i.e., we take the final optimizer learned by MADA and use it as the fixed optimizer from the beginning of training, which we refer to as MADA-FS.
Figure 4: Parameter values with respect to iterations for $\beta_{1,0}=0.9,\beta_{2,0}=0.95,\beta_{3,0}=0$, during training of GPT-2 (125M).
Figure 5: Parameter values with respect to iterations for $\beta_{1,0}=0.7,\beta_{2,0}=0.8,\beta_{3,0}=0$, during training of GPT-2 (125M).
Figure 4: Parameter values with respect to iterations for $\beta_{1,0}=0.9,\beta_{2,0}=0.95,\beta_{3,0}=0$, during training of GPT-2 (125M).
Figure 5: Parameter values with respect to iterations for $\beta_{1,0}=0.7,\beta_{2,0}=0.8,\beta_{3,0}=0$, during training of GPT-2 (125M).
Figure 6: Parameter values with respect to iterations for $\beta_{1,0}=0.9,\beta_{2,0}=0.95,\beta_{3,0}=0.9$, during training of GPT-2 (125M).
5.3 Results
In this section, we first present empirical results of training GPT-2 models on OpenWebText, and provide a discussion of generalization performance and robustness of MADA. We also show how the interpolation coefficients evolve during training. Next, we provide a visualization of the optimal training loss given various initializations of $\beta_{1},\beta_{2}$ values using the smaller Shakespeare dataset. We present additional results in Appendix C including a toy problem to show that MADA discovers solutions that minimize the regret, and comparing MADA and Adam for GPT-2 (medium).
GPT-2 on OpenWebText. When we initialize MADA from AVGrad and Adan with $\beta_{1,0}=0.9,\beta_{2,0}=0.95,\beta_{3,0}=0.9,\rho_{0}=0$ (where subscript 00 denotes the time index) from Figure 3 and Table 3 we observe that MADA consistently outperforms Adam and other recently proposed optimizers, including Lion, Adan, and HyperAdam. In particular, MADA-FS outperforms Adam with established parameters by 0.019 in validation loss which is a significant improvement for this task.
Perplexity on benchmark datasets. To measure the generalization of MADA to other datasets, we compute the validation perplexity of the trained models on three datasets as shown in Table 3. We see that on OpenWebText outperforms Adam and HyperAdam by 0.34 and 0.33 points); while on Wikitext, MADA outperforms these two baselines by 4.45 and 2.37 respectively. On Lambada it is the second-best-performing method with a small gap behind HyperAdam (0.88).
MADA-FS versus MADA. We find that MADA-FS performs slightly better than MADA for OpenWebText (0.004 in validation loss and 0.07, 1.85, 0.68 points on validation perplexity on OpenWebText, Wikitext and Lambada). More details are in Table 3 and Table 3. On the other hand, on Shakespeare we observe that, the dynamically-evolving optimizer performs generally better than using the fixed final optimizer. We conjecture that the difference in the number of iterations or the model size may explain this phenomenon.

### v3

@@ Abstract
Following the introduction of Adam, several novel adaptive optimizers for deep learning have been proposed. These optimizers typically excel in some tasks but may not outperform Adam uniformly across all tasks. In this work, we introduce Meta-Adaptive Optimizers (MADA), a unified optimizer framework that can generalize several known optimizers and dynamically learn the most suitable one during training. The key idea in MADA is to parameterize the space of optimizers and dynamically search through it using hyper-gradient descent during training. We empirically compare MADA to other popular optimizers on vision and language tasks, and find that MADA consistently outperforms Adam and other popular optimizers, and is robust against sub-optimally tuned hyper-parameters. MADA achieves a greater validation performance improvement over Adam compared to other popular optimizers during GPT-2 training and fine-tuning. We also propose AVGrad, a modification of AMSGrad that replaces the maximum operator with averaging, which is more suitable for hyper-gradient optimization. Finally, we provide a convergence analysis to show that parameterized interpolations of optimizers can improve their error bounds (up to constants), hinting at an advantage for meta-optimizers.
@@ 1 Introduction
The choice of an optimization algorithm plays a critical role in determining the downstream performance of a machine learning model. Adaptive moment optimizers are the most preferred class of optimizers employed in most learning tasks such as training Large Language Models (LLMs) (Brown et al., 2020; Touvron et al., 2023) and Diffusion Models (Rombach et al., 2022). In particular, Adam (Kingma and Ba, 2015) is still the “go-to” optimizer in LLM training, despite the emergence of many other optimizers since then.
@@
In this work, we introduce the concept of a parameterized optimizer, which can be viewed as a unified parameterization of a collection of given optimizers. The parameters of a parameterized optimizer define a convex polytope (e.g. hypercube), whose vertices may correspond to individual base optimizers, while the interior represents new optimizers formed by interpolating between them. To make this concept operational, we propose the meta-adaptive optimizer (MADA), which combines the parameterized optimizer with hyper-gradient descent (Baydin et al., 2018) to learn the optimizer coefficients. MADA dynamically adjusts the parameterized optimizer coefficients during training, and effectively adapts the optimizer choice to the learning task. While MADA bears some connections to optimizer search methods, such as (Chen et al., 2023), it does not solely output a final optimizer state. It dynamically adapts the optimizer to the current neighborhood of the loss landscape on-the-fly, removing the need for an offline optimizer search stage, or an outer hyper-parameter selection loop.
Contributions. We make the following contributions:
  • We introduce the concept of a parameterized optimizer, which takes a collection of existing optimizers, and unifies them into a single optimizer, combining the individual update rules through learnable coefficients.
@@
  • We find that not all optimizers are suitable to use in a hyper-gradient optimization framework. Specifically, among popular optimizers, using AMSGrad (Reddi et al., 2018) results in poor performance within MADA due to its use of the maximum operator. Motivated by this, we propose a modification of it called AVGrad, which replaces the maximum operator on the second-order moments with time-averaging, and leads to better performance when used as part of MADA. We also analyze its convergence properties (see Appendix B). We show that MADA converges to AVGrad in a simple example where it is known to perform better than Adam, providing evidence for the effectiveness of MADA in adapting the optimizer to the task (see Appendix C).
  • To demonstrate that optimizer interpolations can improve convergence bounds in an analytically tractable scenario, we theoretically analyze the convergence behavior for interpolations between two optimizers, namely AVGrad and Adam. Our analysis shows that the interpolated optimizer improves the convergence bounds of the base optimizers up to constant factors.
@@
  • We develop a specific parameterized optimizer, which interpolates between Adam (Kingma and Ba, 2015), AVGrad, Yogi (Zaheer et al., 2018), Adan (Xie et al., 2023), and Lion (Chen et al., 2023). On language tasks we compare MADA, which is based on this parameterized optimizer, against Adam, Lion, Adan, and HyperAdam [^1] Throughout this paper we will refer to the version of Adam that uses hyper-gradients to tune $\beta_{1}$ and $\beta_{2}$ parameters, as in Chandra et al. (2022), as HyperAdam., and show that MADA consistently outperforms all baselines. We also illustrate the robustness of MADA to initial hyper-parameters and analyze the evolution of hyper-parameters during training. On vision tasks we compare MADA to Adam, SGD with momentum, and HyperAdam, and observe consistent performance improvement.
@@
Related work. Motivated by the high cost of training large language models, a large number of novel optimizers have been proposed in recent years to speed up the training, increase generalization performance or train more resource-efficient models. As mentioned before, Adam (Kingma and Ba, 2015) is the most commonly employed optimizer, and succeeding methods, in general, try to improve upon it under different scenarios. Reddi et al. (2018); Zaheer et al. (2018) propose AMSGrad and YOGI, respectively, to fix Adam’s potentially non-decreasing effective learning rate. Xie et al. (2023); Dozat (2016) introduce Adan and Nadam, respectively, to replace heavy-ball momentum in Adam with Nesterov momentum. You et al. (2017, 2020) propose LARS and LAMB to improve the performance of Adam in large-batch regime. Heo et al. (2021) proposes AdamP to avoid premature decay of scale-invariant weights. AdaBound (Luo et al., 2019) and AdaBelief (Zhuang et al., 2020) try to estimate a more stable second-order moment term. Foret et al. (2021) proposes sharpness-aware minimization (SAM) and Chen et al. (2020) proposes Padam to increase the generalization performance of the trained models. Liu et al. (2020) stabilizes the training by reducing gradient variance. The work in (Chen et al., 2023) is similar in spirit to our work, in that it introduces a method to symbolically search a space of optimizers; however unlike our work they consider an offline search method, whereas our method learns the optimizer during actual model training, and uses hyper-gradients.
A separate line of work proposed learned optimizers to delegate the optimization task to neural networks (fully connected or LSTMs) (Almeida et al., 2021; Metz et al., 2022, 2020). Instead of directly using first order information as in optimizers, these methods treat gradient information, alongside other features such as training loss, validation loss (Almeida et al., 2021) and so on, as inputs to a neural network which outputs the update to be applied on a particular weight. Learned optimizers require resource-intensive offline training (e.g. thousands of TPU-months in (Metz et al., 2022)) since the updates are handled through a separate neural network that needs to be trained before deployment. Integration of a separate neural network to the optimization process introduces additional complexity, which MADA avoids by requiring only a few parameters to be learned on-the-fly.
@@ 3 Meta-Adaptive Optimizers
4 On Convergence of Interpolated Optimizers
Since MADA uses hyper-gradients to update the optimizer, a prerequisite for a base optimizer to work well within MADA is that its update rules must allow efficient flow of hyper-gradients to the parameterization coefficients. In particular, in our experiments with MADA we have found that including AMSGrad among the base optimizers has an adverse effect on the end performance of the trained model. We conjecture that this is caused by the maximum operator in the second-moment term. Specifically, note that back-propagation of gradients through $\max(a,b)$ corresponds to a simple routing operation to the larger one of $a$ and $b$, where the smaller one is passed 00 gradients. In AMSGrad, $v_{t}=\max\left\{v_{t-1},\bar{v}_{t}\right\}$ which means that for most steps the first term will be greater, causing insufficient hyper-gradient updates on $\beta_{2}$ parameter through $\bar{v}_{t}$ path [^4] To be accurate, considering the example parameterization (3), $\tilde{v}_{t}$ term provides another path for hyper-gradients on $\beta_{2}$. However, in practice we observed that when AMSGrad is used, $\rho$ parameter tends to vanish, diminishing the effect of this path as well.. To remedy this, we introduce AVGrad (formally defined in next section), which replaces the maximum operator with time-averaging of second moments, and results in better hyper-gradient flow and validations loss.
4.1 AVGrad and its interpolation with Adam
@@
where $\tilde{v}_{t}$ is the running average of past $\overline{v}_{t}$’s. We provide the convergence analysis of AVGrad in Appendix B in Theorems 2 and 3
In what follows, we focus on the interpolation between Adam and AVGrad, with the interpolation coefficient $\rho_{t}$. Specifically, we replace last line in (5) with $v_{t}:=\rho_{t}\bar{v}_{t}+(1-\rho_{t})\tilde{v}_{t}$ where $\rho_{t}$ interpolates between second-order moments of Adam and AVGrad.
@@
In Appendix C, we use MADA to solve the convex problem given in (Reddi et al., 2018), an example where Adam fails. We found that MADA quickly converges to AVGrad, and thus to the optimum, by adapting $\rho_{t}\rightarrow 0$, even when we initialize it from Adam ($\rho_{0}=1$).
@@ 4.2 Convergence of the interpolated optimizer
Due to the complexity of our overall parameterization, deriving theoretical guarantees for MADA at its full scope is challenging. Therefore, in this section, we reduce the scope by examining convergence properties of interpolations between two optimizers, namely, AVGrad and Adam. In this setting, we will show that the parameterized optimizer allows the design of novel interpolated optimizers which improve the convergence rate of either optimizer up to constant factors, demonstrating the value of the parameterized optimizer formulation. We defer the proof and other technical details to Appendix A.
We make the following standard assumptions: $F$ is bounded from below, is $L$-smooth, and with stochastic gradients that are bounded almost surely.
Défossez et al. (2022) proposed a unified analysis for the convergence analysis of Adagrad and Adam in the non-convex setting. Since AVGrad can be seen as a version of Adagrad, the proof technique of Défossez et al. (2022) can be adapted to AVGrad with some modifications. The modification includes the introduction of $\rho$ parameter which interpolates between the second-order moment and the moment average, and results in an additional degree of freedom. Subsequently, $\rho$ can be chosen to skew the updates towards the advantageous term under different scenarios.
Here we state the convergence result of the interpolated optimizer without momentum, that is when $\beta_{1}=0$ (see precise statement in Appendix A). The extension to the case with momentum can be done in the same way we extend Theorem 2 to Theorem 3, similar to (Défossez et al., 2022).
@@
Remark. Aligned with the condition in Proposition 1, the contribution from $v_{t}$ in Theorem 1 is multiplied with $1/t$ in order to avoid divergent terms in the bound. Observe that when $\rho=0$ or $\rho=1$ we recover the convergence rates for AVGrad (see Theorem 2 in Appendix B) and Adam (Défossez et al., 2022) respectively. We note that for the right-hand side to remain bounded, one would need either $0\leq\rho\leq\beta_{2}$ (all terms vanish as in AVGrad), or $\rho=1$ (a constant term remains as in Adam). To gain more insights, in Figures 2 and 2, we plot the value of the bound in Theorem 1 with respect to $\rho$ for two representative cases . The first case with larger $T$ represents a more favorable setting for AVGrad as all the terms vanish with $T$. The second is more favorable for Adam since the vanishing terms vanish faster than AVGrad. In both cases, the interpolated optimizer provides the best convergence bound when $\rho=\beta_{2}$, which suggests an advantage for the interpolated optimizers.
@@ 5 Experiments
In our experiments, we aim to answer how the generalization and downstream performance of MADA compares against fixed optimizers, and how robust MADA is to poor hyper-parameter initializations compared to fixed optimizers. We focus on the pre-training of models from scratch. We also try to gain insights into the behavior of MADA by monitoring the evolution of the optimizer coefficients. Additional details of our experimental setup, and results can be found in Appendix C.
Figure 3: Validation losses of competing methods on OpenWebText for GPT-2 (125M) model using the same random seed.
Figure 4: Training losses of competing methods on OpenWebText for GPT-2 (125M) model using same random seed, smoothed for convenience.
@@
5.1 A concrete parameterized optimizer
5.2 Experimental setting
Data and models. In recent years, auto-regressive models have been widely used for benchmarking and algorithm evaluation (Radford et al., 2019b; Liu et al., 2023). Motivated by this, we evaluate MADA on the causal language modeling task with GPT-2, over two datasets: Shakespeare (Karpathy, 2015), and OpenWebText (Gokaslan and Cohen, 2019). On the Shakespeare dataset, we train a 6 layer transformer with context size of 256 (11M parameters) and fine-tune GPT-2 (XL) with 48 layers (1.5B parameters). On OpenWebText, we train GPT-2 (small) with 12 layers, 1024 context size (125M parameters) and GPT-2 (medium) with 24 layers, 1024 context size (335M parameters).  [^5] We use nanoGPT (https://github.com/karpathy/nanoGPT) code base for the implementation.
Baselines and learned optimizers. On OpenWebText, we compare MADA to Adam (Kingma and Ba, 2015), HyperAdam (Chandra et al., 2022) (Adam with $\beta_{1}$ and $\beta_{2}$ parameters learned through hyper-gradients), Adan (Xie et al., 2023), Lion (Chen et al., 2023), and AVGrad. For all the methods we used decoupled weight decay (Loshchilov and Hutter, 2019), i.e. the AdamW variant of Adam is used. We keep the learning hyper-parameters such as learning rate schedule and weight decay identical, except for Lion where we choose a lower initial learning rate as suggested in (Liu et al., 2023). For a fair comparison, for all optimizers, we use the same codebase based on Chandra et al. (2022). For OpenWebText experiments, we use established parameters for Adam ($\beta_{1}=0.9,\beta_{2}=0.95,\epsilon=10^{-6}$) and also use these values as the initial parameters for MADA, HyperAdam, and AVGrad; for other methods we use the parameters suggested in respective papers; and measure the validation loss. For Shakespeare experiments, we compare MADA to Adam and HyperAdam; the relatively small model size in this experiment allows us to sweep the grid of initial $\beta$ parameters and compare the final training loss across many different hyper-parameter choices. In some experiments, we also evaluate the final state of MADA statically, i.e., we take the final optimizer learned by MADA and use it as the fixed optimizer from the beginning of training, which we refer to as MADA-FS.
Figure 5: Parameter evolution for $\beta_{1,0}=0.9,\beta_{2,0}=0.95,\beta_{3,0}=0$.
Figure 6: Parameter evolution for $\beta_{1,0}=0.7,\beta_{2,0}=0.8,\beta_{3,0}=0$.
Figure 7: Parameter evolution for $\beta_{1,0}=0.9,\beta_{2,0}=0.95,\beta_{3,0}=0.9$.
Tuning hyper-learning rates. While hyper-learning rates are additional hyper-parameters to be tuned; the tasks are less sensitive to hyper-learning rates compared to other hyper-parameters such as learning rate, most likely because each hyper-learning rate controls the update of a single parameter. As a result, hyper-learning rates can be fine-tuned easily and can be transferred across similar tasks. In general, for the hyper-learning rates of $\beta_{1},\beta_{2}$, since the gradients are relatively large, we search in a set of smaller values: $\{10^{-3},5\times 10^{-4},10^{-4}\}$. For the other parameters we search in the set $\{10^{-1},5\times 10^{-2},10^{-2}\}$.
5.3 Results
In this section, we first present empirical results of training GPT-2 models on OpenWebText, and provide a discussion of generalization performance and robustness of MADA. We also show how the interpolation coefficients evolve during training. Next, we provide a visualization of the optimal training loss given various initializations of $\beta_{1},\beta_{2}$ values using the smaller Shakespeare dataset. We present additional results in Appendix C including a toy problem to show that MADA discovers optimal solutions that Adam cannot find, and comparing MADA and Adam for GPT-2 (medium).
GPT-2 on OpenWebText. When we initialize MADA from AVGrad and Adan with $\beta_{1,0}=0.9,\beta_{2,0}=0.95,\beta_{3,0}=0.9,\rho_{0}=0$ (where subscript 00 denotes the time index) from Table 2 we observe that MADA consistently outperforms Adam and other recently proposed optimizers, including Lion, Adan, and HyperAdam. In particular, MADA-FS outperforms Adam with established parameters by 0.019 in validation loss which is a significant improvement for this task. In Figure 4 and Figure 4 we see that MADA is able to converge to a lower loss than baseline methods (both in terms of validation and training loss). Note that MADA is able to converge faster than AVGrad and Adam with the same learning rate and schedule, which may be attributed to theoretical result Theorem 1 on interpolated optimizers.
Perplexity on benchmark datasets. To measure the generalization of MADA to other datasets, we compute the validation perplexity of the trained models on three datasets as shown in Table 3. We see that on OpenWebText MADA outperforms Adam and HyperAdam by 0.34 and 0.33 points); while on Wikitext (Merity et al., 2016), MADA outperforms these two baselines by 4.45 and 2.37 respectively. On Lambada (Radford et al., 2019a) it is the second-best-performing method with a small gap behind HyperAdam (0.88).
MADA-FS versus MADA. We find that MADA-FS performs slightly better than MADA for OpenWebText (0.004 in validation loss and 0.07, 1.85, 0.68 points on validation perplexity on OpenWebText, Wikitext and Lambada). More details are in Tables 2 and 3. On the other hand, on Shakespeare we observe that, the dynamically-evolving optimizer performs generally better than using the fixed final optimizer. We conjecture that the difference in the number of iterations, the model size or dataset may explain this phenomenon.